A train leaves Kansas City for a nearby city at 8:15 pm, averaging 90 mph.
Another train leaves the nearby city for Kansas City at 10:45 pm, averaging 120
mph.
If the nearby city is 4215 miles from Kansas City, to the nearest minute, at what
time will the two trains pass each other?

Found 3 solutions by greenestamps, ewatrrr, ikleyn:
Answer by greenestamps(13200)



The first train travels for 2:30 at 90 mph before the second train starts, traveling 90(2.5) = 225 miles in that time. The remaining distance between the two trains is 4215-225 = 3990 miles.

When the two trains are both moving, they are closing the distance between them at a rate of 90+120 = 210 mph.

The number of hours required for them to travel the 3990 miles that remain before they meet is 3990/210 = 19.

ANSWER: The two trains will pass each other 19 hours (5 hours less than a full day) after the second train started, at 5:45 pm the next day.

--------------------------------------------------

NOTES for tutor ewatrrr....

(1) The question you answered is not at all the question that was asked....
(2) please turn off your special font type and color at the end of your responses so it doesn't carry over to the responses from other tutors


Answer by ewatrrr(24785)

 
Hi,
Let t represent time nearby town train travels prior to meeting
Note: 10:45 - 8:15 = 2.5 hrs additional time time KC train travels
90(t+ 2.5) + 120t = 4215
210t = 3990
t = 19 hr time nearby town train travels prior to meeting
Left at 10:45 pm
19hr later... 1:15 to midnite + 17 3/4 = 19hr
Trains meet at 5:45pm the next day
Wish You the Best in your Studies.

Answer by ikleyn(52792)

.

Do you know the meaning of the English word "nearby" ?


Consult in any dictionary . . . and use words properly . . .


Otherwise,  people will laugh reading your "problems" . . .

while other people may feel themselves very bad reading it, and will require urgent medical help . . .


I personally belong to the first category, so thanks for making me smiling . . .


-------------

Hello, this day I saw so many nonsensical posts,  that I decided to make this announcement:


        If I see tomorrow nonsensical posts,  I will collect all relevant links to them.

        After that,  I will post these links to the management of this project,  asking
        to remove their authors from their positions due to absolute absence of skills
        to formulate their posts correctly.